# Continuum Attention for Neural Operators

### v3

@@ 6.2.2 Darcy Flow
For the low resolution setting, we train the FNO, AFNO, Galerkin transformer and TNO architectures using $3600$ independent samples of resolution $64\times 64$, we use $200$ samples for validation and $200$ samples for testing. The Fourier neural operator is parametrized by $12$ Fourier modes in each dimension and channel width of $128$; the model is trained using a batch size of $8$ and learning rate of $10^{-4}$. On the other hand, the transformer neural operator is parametrized by $128$ channels in the encoder and is trained using a batch size of 22 and learning rate of $10^{-3}$. The AFNO is instantiated with $4$ layers of $128$ hidden channels, and does not employ patching; it is trained using a learning rate of $10^{-3}$ and batch size of $1$. The Galerkin transformer uses the Fourier type attention from Cao (2021), with quadratic complexity, making it prohibitive for higher resolutions. It uses $4$ layers of $128$ hidden channels and is trained using a learning rate of $10^{-3}$ and batch size of $2$. On this problem we train all neural operators using the relative $H^{1}$ loss.
@@ 6.2.3 Kolmogorov Flow
The dot on $L^{2}$ denotes the space of spatially mean-zero functions. The vorticity is defined as $w=(\nabla\times u)\widehat{z}$ and the stream function $f$ as the solution to the Poisson equation $\Delta f=w$. Existence of the semigroup $S_{t}:\mathcal{U}\to\mathcal{U}$ is shown in Temam (2012, Theorem 2.1). We generate the data by solving (KF) in vorticity-streamfunction form by applying the pseudo-spectral split step method from Chandler and Kerswell (2013). In our experimental set-up $n=4$ and $\nu=1/70$ are chosen. Random initial conditions are sampled from the Gaussian measure $\mathcal{N}(0,C)$ where the covariance operator is given by
@@ 7 Conclusions
In this work we have introduced a continuum formulation of the attention mechanism from the seminal work Vaswani et al. (2017). Our continuum formulation can be used to design transformer neural operators. Indeed, this continuum formulation leads to discretization invariant implementations of attention and schemes that exhibit zero-shot generalization to different resolutions. In the first result of its kind for transformers, with a slight modification to the architecture implemented in practice, the resulting neural operator architecture mapping between infinite-dimensional spaces of functions is shown to be a universal approximator of continuous functions and functions of Sobolev regularity defined over a compact domain. We extend the continuum formulation to patched attention, which makes it possible to design efficient transformer neural operators. To this end, we introduce a neural operator analogue of the vision transformer Dosovitskiy et al. (2021) and a more expressive architecture, the Fourier attention neural operator. Through a cost-accuracy analysis, we demonstrate the power of the methodology in the context of a range of operator learning problems. In the following we highlight potential avenues for further work.
@@
    The Fourier attention neural operator architecture uses operator parametrizations for $\mathsf{Q},\mathsf{K},\mathsf{V}$. Other design choices for these operators not employing the Fourier basis are possible, for example using the wavelet basis as in Tripura and Chakraborty (2023) or using a multigrid approach as in He et al. (2024). An investigation of other attention neural operator architectures arising from this consideration would be of interest.
@@ Appendix C Complexity of the Transformer Neural Operators
We summarize the expressions for the parameter complexity and evaluation cost (measured in floating point operations, FLOPS) of the neural operators from Section 4 and the Fourier neural operator (FNO) from Li et al. (2021). These are provided in Tables 4 and 5, respectively. In Subsection C.1 and C.2 we provide further details for the computations of the parameter and evaluation cost complexities, respectively.
@@ C.1 Parameter Complexity
Table 4: Parameter complexity for the three proposed transformer neural operators, as implemented in practice, compared to the Fourier neural operator (Li et al., 2021). We note that for our implementations, $d_{\textrm{FNO}}$, which is the latent dimension of the two-layer lifting and projection MLP in FNO, is fixed to be $128$, while the hyperparameter $b$ in the AFNO is fixed to be $16$. Parameters arising from possible bias terms, parameters arising from skip connections in FNO and parameters arising from normalizations in AFNO are omitted for ease of presentation.
@@ C.2 Evaluation Cost
We note that a matrix vector product $Wu$ for $W\in\mathbbm{R}^{r\times r^{\prime}}$ has cost $2rr^{\prime}$. We approximate the cost of computing the FFT of a vector in $\mathbbm{R}^{r}$ as $5r\log r$ (Hoop et al., 2022). We approximate the cost of the two-dimensional real FFT, used in implementation, by $(15/2)\cdot N\log(\sqrt{N})$ where $N$ is the total number of grid points. Assuming that the grid is of size $s\times s=N$ for $s$ an even integer, 1d FFTs are first computed for each row. The real FFT requires $s/2$ column 1d FFTs, hence the $(15/2)\cdot N\log\sqrt{N}$ complexity.
@@
Table 5: Evaluation cost (in FLOPS) for the proposed transformer neural operators, as implemented in practice, compared to FNO (Li et al., 2021). As done in Kaplan et al. (2020) and for ease of presentation, the additional evaluation cost arising from nonlinear activations, the softmax operator and normalizations are omitted. Furthermore, we omit the cost arising from skip connections. Indeed these do not affect the scaling. We further note that for our implementations, $d_{\textrm{FNO}}$, which is the latent dimension of the two-layer lifting and projection MLP in FNO, is fixed to be $128$. In this table we present the expression for the evaluation cost of AFNO without consideration of possible patching, which we do not employ for our comparisons: this in combination with the fact that we do not employ mode truncation is the reason for the $Nd_{\textrm{model}}^{2}$ term. In our context, the parameter $b$ in AFNO is set to $b=8$.

### v4

@@ 6.2.2 Darcy Flow
For the low resolution setting, we train the FNO, AFNO, Galerkin transformer and TNO architectures using $3600$ independent samples of resolution $64\times 64$, we use $200$ samples for validation and $200$ samples for testing. The Fourier neural operator is parametrized by $12$ Fourier modes in each dimension and channel width of $128$; the model is trained using a batch size of $8$ and learning rate of $10^{-4}$. On the other hand, the transformer neural operator is parametrized by $128$ channels in the encoder and is trained using a batch size of 22 and learning rate of $10^{-3}$. The AFNO is instantiated with $4$ layers of $128$ hidden channels, and does not employ patching; it is trained using a learning rate of $10^{-3}$ and batch size of $1$. The Galerkin transformer uses the Fourier type attention from cao2021choose, with quadratic complexity, making it prohibitive for higher resolutions. It uses $4$ layers of $128$ hidden channels and is trained using a learning rate of $10^{-3}$ and batch size of $2$. On this problem we train all neural operators using the relative $H^{1}$ loss.
@@ 6.2.3 Kolmogorov Flow
The dot on $L^{2}$ denotes the space of spatially mean-zero functions. The vorticity is defined as $w=(\nabla\times u)\widehat{z}$ and the stream function $f$ as the solution to the Poisson equation $\Delta f=w$. Existence of the semigroup $S_{t}:\mathcal{U}\to\mathcal{U}$ is shown in temam2012infinite. We generate the data by solving (KF) in vorticity-streamfunction form by applying the pseudo-spectral split step method from Chandler_Kerswell_2013. In our experimental set-up $n=4$ and $\nu=1/70$ are chosen. Random initial conditions are sampled from the Gaussian measure $\mathcal{N}(0,C)$ where the covariance operator is given by
@@ 7 Conclusions
In this work we have introduced a continuum formulation of the attention mechanism from the seminal work vaswani2017attention. Our continuum formulation can be used to design transformer neural operators. Indeed, this continuum formulation leads to discretization invariant implementations of attention and schemes that exhibit zero-shot generalization to different resolutions. In the first result of its kind for transformers, with a slight modification to the architecture implemented in practice, the resulting neural operator architecture mapping between infinite-dimensional spaces of functions is shown to be a universal approximator of continuous functions and functions of Sobolev regularity defined over a compact domain. We extend the continuum formulation to patched attention, which makes it possible to design efficient transformer neural operators. To this end, we introduce a neural operator analogue of the vision transformer dosovitskyi2021image and a more expressive architecture, the Fourier attention neural operator. Through a cost-accuracy analysis, we demonstrate the power of the methodology in the context of a range of operator learning problems. In the following we highlight potential avenues for further work.
@@
    The Fourier attention neural operator architecture uses operator parametrizations for $\mathsf{Q},\mathsf{K},\mathsf{V}$. Other design choices for these operators not employing the Fourier basis are possible, for example using the wavelet basis as in TRIPURA2023115783 or using a multigrid approach as in he2024mgno. An investigation of other attention neural operator architectures arising from this consideration would be of interest.
@@ Appendix C Complexity of the Transformer Neural Operators
We summarize the expressions for the parameter complexity and evaluation cost (measured in floating point operations, FLOPS) of the neural operators from Section 4 and the Fourier neural operator (FNO) from li2021fourier. These are provided in Tables 4 and 5, respectively. In Subsection C.1 and C.2 we provide further details for the computations of the parameter and evaluation cost complexities, respectively.
@@ C.1 Parameter Complexity
Table 4: Parameter complexity for the three proposed transformer neural operators, as implemented in practice, compared to the Fourier neural operator (li2021fourier). We note that for our implementations, $d_{\textrm{FNO}}$, which is the latent dimension of the two-layer lifting and projection MLP in FNO, is fixed to be $128$, while the hyperparameter $b$ in the AFNO is fixed to be 88. Parameters arising from possible bias terms, parameters arising from skip connections in FNO and parameters arising from normalizations in AFNO are omitted for ease of presentation.
@@ C.2 Evaluation Cost
We note that a matrix vector product $Wu$ for $W\in\mathbbm{R}^{r\times r^{\prime}}$ has cost $2rr^{\prime}$. We approximate the cost of computing the FFT of a vector in $\mathbbm{R}^{r}$ as $5r\log r$ (hoop2022cost). We approximate the cost of the two-dimensional real FFT, used in implementation, by $(15/2)\cdot N\log(\sqrt{N})$ where $N$ is the total number of grid points. Assuming that the grid is of size $s\times s=N$ for $s$ an even integer, 1d FFTs are first computed for each row. The real FFT requires $s/2$ column 1d FFTs, hence the $(15/2)\cdot N\log\sqrt{N}$ complexity.
@@
Table 5: Evaluation cost (in FLOPS) for the proposed transformer neural operators, as implemented in practice, compared to FNO (li2021fourier). As done in Kaplan2020ScalingLF and for ease of presentation, the additional evaluation cost arising from nonlinear activations, the softmax operator and normalizations are omitted. Furthermore, we omit the cost arising from skip connections. Indeed these do not affect the scaling. We further note that for our implementations, $d_{\textrm{FNO}}$, which is the latent dimension of the two-layer lifting and projection MLP in FNO, is fixed to be $128$. In this table we present the expression for the evaluation cost of AFNO without consideration of possible patching, which we do not employ for our comparisons: this in combination with the fact that we do not employ mode truncation is the reason for the $Nd_{\textrm{model}}^{2}$ term. In our context, the parameter $b$ in AFNO is set to $b=8$.